# WeatherEdit: Controllable Weather Editing with 4D Gaussian Field

### v2

@@ Abstract
In this work, we present WeatherEdit, a novel weather editing pipeline for generating realistic weather effects with controllable types and severity in 3D scenes. Our approach is structured into two key components: weather background editing and weather particle construction. For weather background editing, we introduce an all-in-one adapter that integrates multiple weather styles into a single pretrained diffusion model, enabling the generation of diverse weather effects in 2D image backgrounds. During inference, we design a Temporal-View (TV-) attention mechanism that follows a specific order to aggregate temporal and spatial information, ensuring consistent editing across multi-frame and multi-view images. To construct the weather particles, we first reconstruct a 3D scene using the edited images and then introduce a dynamic 4D Gaussian field to generate snowflakes, raindrops and fog in the scene. The attributes and dynamics of these particles are precisely controlled through physical-based modelling and simulation, ensuring realistic weather representation and flexible severity adjustments. Finally, we integrate the 4D Gaussian field with the 3D scene to render consistent and highly realistic weather effects. Experiments on multiple driving datasets demonstrate that WeatherEdit can generate diverse weather effects with controllable condition severity, highlighting its potential for autonomous driving simulation in adverse weather. See project page: https://jumponthemoon.github.io/w-edit
{strip}[Uncaptioned image]
Figure 1: Realistic, Controllable Weather Editing. Given input sunny weather images from multi-view cameras across multiple timesteps, WeatherEdit  can generate high-fidelity scenarios under diverse weather conditions (e.g., foggy, snowy, and rainy) with various severity levels (e.g., light, moderate, heavy) based on user instructions. The best view is in zoom.
1 Introduction
@@
3D weather editing (3D-WE) aims to generate multi-weather conditions in 3D scenes based on user instructions. It is a crucial task in computer vision with applications in autonomous driving [27, 26], augmented reality, and virtual scene synthesis [1]. While existing methods range from 2D-based image techniques [36, 41] to 3D neural rendering [17], developing a unified framework that enables multiple realistic weather effects and controllable severity remains a significant challenge. For instance, style transfer methods can modify weather backgrounds but are typically limited to a single effect. Diffusion models [26, 39] enable multi-weather synthesis; however, they often lack fine-grained control and struggle with spatial and temporal consistency across multiple images [30]. More recently, ClimateNeRF [17] utilized Neural Radiance Fields (NeRF) to model snow accumulation and flooding in 3D space. However, this method is limited to static weather events and leaves the overall weather tone unchanged, lacking full-scene atmospheric realism.
@@
    We demonstrate WeatherEdit’s effectiveness in generating realistic, consistent, and controllable weather effects in 3D driving scenes, showcasing its applicability to real-world scenarios.
2 Related Work
@@
Existing methods can be categorized into 2D-based and 3D-based approaches. 2D-based methods utilize generative models [43, 24] to synthesize weather effects in single images. [14, 33] use CycleGAN [43] to edit weather conditions, while [32] applies GAN-based inpainting to generate flooding with water reflections. [11] uses graphical model to produce fog effects, while [26] finetunes a diffusion model to various weather effects. Conventional 3D-based methods [34, 44, 8, 10] use graphical simulation to model weather in 3D space. More recently, [17] has introduced neural rendering techniques for creating weather backgrounds but it is limited to post-weather event.
@@
Diffusion models [12, 21, 29] have significantly advanced image editing, enabling flexible and controllable modifications. Methods such as ControlNet [42] and T2I-Adapter [23] introduce trainable modules to guide the generation process, allowing for fine-grained control over diffusion-based outputs. InstructPix2Pix [4] enables interactive editing with textual instructions by trained on image-text pairs. TurboEdit [7] employs a pseudo-guidance approach to enhance edit magnitude without introducing artifacts.
@@
Neural Radiance Fields (NeRF) [22]and 3D Gaussian Splatting (3DGS) [15] are leading techniques for 3D scene reconstruction from multi-view images. NeRF encodes scenes implicitly with a neural network, while 3DGS represents them explicitly as 3D Gaussians, enabling faster rendering. Their versatility allows applications in dynamic scene reconstruction [6, 38], content creation [20, 19], and artifact removal [25]. While [17] extends NeRF to model smog, snow, and floods for weather simulation but focuses only on static weather conditions, leaving dynamic weather phenomena unaddressed.
3 Methods
In order to generate realistic and controllable multi-weather scenes in 3D, we propose a novel framework, WeatherEdit, which follows a progressive 2D-to-4D transformation process. As illustrated in Fig. 2, WeatherEdit  adopts a two-step weather editing pipeline: weather background editing and weather particle construction. The first step generates distortion-free, spatially and temporally aligned 2D images with edited backgrounds, while the second step dynamically models weather particles with various types and severity levels. The joint association of these two steps enables the flexible generation of multi-weather scenes.
3.1 Weather Background Editing
All-in-one Adapter. Existing diffusion-based methods often suffer from scene deformations and excessive stylization, making them unsuitable for precise and realistic weather editing. While [24] achieves effective weather editing by fine-tuning separate models with multiple LoRA [13] adapters (shown in Fig. 3 (a)) for different styles, it requires multiple retraining steps, leading to inefficiencies in storage and training time.
Figure 3: Single Adapter vs. All-in-One Adapter. (a) Utilizes separate adapters for each weather style (e.g., snow, rain, fog), while (b) employs a unified all-in-one adapter for efficient multi-style adaptation. Blue box highlights adapters to be trained.
To address this, we propose an all-in-one LoRA adapter that injects multiple weather styles into a single model (as shown in Fig. 3 (b),), eliminating the need for training multiple ones. Let $W_{0}\in\mathbb{R}^{d\times k}$ be the pre-trained weights, $x$ the input image, and $i$ the weather style (e.g., snowy, rainy, foggy) specified by text prompt. Each style is associated with a low-rank matrix $L_{i}$. The adapted model’s forward pass is then computed as $h=W_{0}x+L_{i}x$.
@@ 3.1 Weather Background Editing
Temporal-View Consistency Allignment.
@@
After fine-tuning the model, multi-time and multi-view images are processed to achieve background scene editing. However, processing each image independently can lead to inconsistencies across time steps and viewpoints due to the inherent stochasticity of generative models. To address this, we propose a Temporal-View (TV-) attention mechanism to enable our model to integrate information from adjacent frames and key viewpoints.
@@
[…]
@@
To preserve temporal coherence, each frame $t$ queries its adjacent frames ($t-1$ and $t+1$), allowing it to integrate contextual information from both past and future frames. The temporal attention is defined as:
[…]
@@
3.2 Weather Particle Construction
Real-world weather is inherently regional, and we incorporate this into weather particle construction. Specifically, we propose a 4D Gaussian field, initialized with 3D Gaussians carrying physical attributes (e.g., color, quantity) to model raindrops, fog, and snowflakes through attribute modeling. To enhance realism, we apply dynamic simulation to create realistic weather effects.
@@ 3.2 Weather Particle Construction
Figure 6: Comparison with 2D-based Editing Methods. We show (a) original images with weather editing results from (b) ControlNet, (c) InstructPix2Pix, (d) TurboEdit, and (e) Ours. Existing methods often suffer from over-stylization, content removal, and inconsistencies in spatial and temporal coherence. In contrast, our method preserves scene integrity while ensuring multi-view and multi-frame consistency, producing realistic and temporally coherent weather effects across viewpoints and frames.
@@ 4.1 Experimental Setup
Dataset. To develop the weather dataset for all-in-one adapter fine-tuning, we extract 1,237 image pairs from the BDD100K [40], MUSE [3], and ACDC [31] datasets. Each pair contains a normal-weather image and an adverse-weather image, including snowy, rainy, and foggy conditions, along with a corresponding ground-truth semantic segmentation map. For weather editing, we selected eight scenes with multi-view image sequences from the Pandaset [37], Waymo Open Dataset [35], nuScenes [5], and KITTI-360 [18] datasets.
@@
Implementation details. We implement the 2D editing model using CycleGAN-Turbo[24] and train on a single RTX A6000. The 4D Gaussian field and 3D scene reconstruction is based on OmniRe[6] and 3DGS[15], with both training and rendering performed on a single RTX 3090.
@@
Evaluation Metrics. To evaluate edited weather background quality, we use the cosine similarity of CLIP [28] image embeddings (CLIP-S) to measure content preservation and the directional CLIP similarity [9] (CLIP-DS) to assess alignment with text instructions. To assess consistency across temporal and spatial dimensions, we calculate the warp error to measure temporal coherence between adjacent frames and employ the Bhattacharyya distance [2] to quantify color distribution similarity across viewpoints.
@@
Figure 8: Controllable Weather Effect and Severity in Driving Scenes. WeatherEdit can modify (a) original driving scenes from Pandaset, Nuscenes, and Waymo datasets by generating (b) snowy, foggy, and rainy effects while preserving scene structure. Additionally, it enables (c) controllable weather severity, adjusting precipitation intensity from light to heavy for realistic and flexible weather simulation.
4.2 Experimental Results
@@
We present both qualitative and quantitative comparisons of our weather background editing method against 2D-based approaches, including ControlNet [42], InstructPix2Pix [4], and TurboEdit [7], as well as 3D-based methods, such as 3D stylization [16] and ClimateNeRF [17]. For dynamic weather effects, we apply qualitative analysis, i.e., controllable weather simulation, to demonstrate the results.
@@
Weather Background Editing Comparison. For the comparison of 2D-based methods, as shown in Fig. 6, we observe that ControlNet tends to over-stylize the image, altering its original content. InstructPix2Pix produces a reasonable effect; however, the effect is inconsistent across adjacent frames and viewpoints. While TurboEdit accurately stylizes the image, it significantly modifies the scene content (e.g., turning pedestrians into water reflections and trees in a foggy scene). In contrast, our method achieves a realistic and coherent weather transformation while preserving scene integrity, as also indicated by the highest CLIP-DS and CLIP-S scores shown in Tab. 1.
ControlNet InstructPix2Pix TurboEdit Ours
CLIP-DS Score ↑ 0.2905 0.2953 0.2573 0.3023
CLIP-S Score ↑ 0.5944 0.6908 0.6762 0.7495
Table 1: Comparison with 2D-based weather background editing.
@@
CLIP-DS Score ↑ 0.2253 0.2674 0.2834
CLIP-S Score ↑ 0.8417 0.7022 0.7513
Table 2: Comparison with 3D-based weather background editing.
For 3D-based methods, 3D stylization produces only a barely perceptible snowy effect on the vegetation, aligning with the highest CLIP-S score(shown in Tab. 2), which indicates minimal change in the scene. ClimateNeRF fails to modify the overall scene tone to accurately reflect the intended weather conditions. In contrast, our approach not only synthesizes a realistic snowy effect but also adjusts the scene tone to create a coherent and perceptually convincing snowy atmosphere, as also indicated by the highest CLIP-DS score in Tab. 2.
Controllable Weather Simulation. A key advantage of our method over existing approaches is its ability to generate dynamic weather effects while precisely controlling both the type and severity of weather conditions. Additionally, we can flexibly choose whether to introduce weather particles and, if included, adjust their intensity with fine granularity. This allows for a range of scenarios, from a static wet road surface without active rainfall to a heavy snowstorm with dense snowfall obscuring visibility. To visualize the results, we render edited weather backgrounds and modulate weather severity. The weather backgrounds, depicted in Fig. 8b, are reconstructed from 2D editing outputs, with the specific weather type determined by the selected adapters during the editing phase. Following this, weather particles are introduced by applying the 4D Gaussian field, as illustrated in Figure Fig. 8c. By adjusting particle attributes such as color and opacity, we can simulate various weather particles, including snowflakes, raindrops and fog. Moreover, by controlling particle quantity, shape, rotation, and velocity, we can achieve different levels of severity—ranging from light to moderate to heavy precipitation. Our approach enables highly flexible and physics-consistent scene manipulation, allowing for the realistic simulation of diverse environmental conditions.
5 Ablation Studies
@@
Semantic Conditioned Input. We investigate the effectiveness of semantic-conditioned input by comparing editing results and FID score with and without it. By computing the cross-attention map using ”snowy” in the text prompt, we observe that semantic conditioning enhances the model’s response to sidewalks and trees, aligning weather effects more accurately with scene structures, as shown in Fig. 11. Additionally, the model achieves a lower average FID value across different weather conditions compared to cases without semantic conditioning, demonstrating its effectiveness in capturing weather characteristics during training.
Temporal-View Attention. As shown in Fig. 10, without TV-Attention, view 3 exhibits inconsistent editing effects, with a darker tone across two frames compared to view 2. After incorporating temporal-view attention, view 3 achieves more consistent editing, displaying a similar snowy effect between adjacent views and frames. For the qualitative analysis shown in Tab. 3, temporal attention reduces warp error, indicating improved temporal coherence, while spatial attention results in a lower Bhattacharyya distance, suggesting more consistent editing across viewpoints. By integrating both temporal and view attention, the full implementation achieves the lowest Bhattacharyya distance while maintaining a balanced reduction in warp error, ensuring globally consistent editing results.
Figure 9: Ablation Study on WeatherEdit modules. (b) Without Weather Background Editing, the scene lacks atmospheric tone. (c) Without Attribute Modeling, the scene presents unrealistic, blobs instead of proper weather effects. (d) Without Dynamic Simulation, particles remain static in the camera view (highlighted in red circle), lacking natural motion. (e) Without Local Field Alignment, weather particles stay fixed at their initial position and disappear as the camera moves forward (highlighted in red line). (f) Full Implementation generates a realistic weather background with natural, scene-consistent weather dynamics.
@@
Self-attention only w View w Temporal Full
Warp-error $\downarrow$ 0.043 0.042 0.039 0.041
Bhattacharyya distance $\downarrow$ 0.272 0.253 0.261 0.245
@@
Modular Design for 4D Gaussian Field. As shown in Fig. 9, without 2D background editing (b), the scene contains weather particles but lacks an overall weather tone. During attribute modeling, we randomly initialize 4D Gaussians, resulting in a scene filled with colorful blobs (c) instead of realistic weather effects. Without dynamic simulation (d), the particles remain relatively static in the camera view, lacking the natural falling motion. Without local field alignment (e), weather particles stay fixed at the initial position of the scene. As the camera moves forward, the weather particles gradually disappear instead of persisting throughout the scene. In contrast, when all modules are included, the full implementation (f) produces a scene with a realistic weather background and natural weather dynamics.
6 Conclusion
We introduce WeatherEdit, a novel 4D weather editing framework capable of generating realistic weather effects—including fog, rain, and snow—with precise control over their severity. Our approach leverages an all-in-one adapter that enables a diffusion model to transform normal-weather images into diverse weather styles. To maintain consistency of edited results across multiple frames and viewpoints, we introduce an ordered Temporal-View (TV) Attention mechanism. Additionally, by designing a 4D Gaussian field, we model the attributes and dynamics of weather particles within a 3D scene, enhancing realism and enabling fine-grained control over weather severity. WeatherEdit demonstrates its potential in generating extreme weather conditions from normal scenes, providing a foundation for future research on evaluating the resilience of real-world applications such as autonomous driving.

### v3

@@ Abstract
In this work, we present WeatherEdit, a novel weather editing pipeline for generating realistic weather effects with controllable types and severity in 3D scenes. Our approach is structured into two key components: weather background editing and weather particle construction. For weather background editing, we introduce an all-in-one adapter that integrates multiple weather styles into a single pretrained diffusion model, enabling the generation of diverse weather effects in 2D image backgrounds. During inference, we design a Temporal-View (TV-) attention mechanism that follows a specific order to aggregate temporal and spatial information, ensuring consistent editing across multi-frame and multi-view images. To construct the weather particles, we first reconstruct a 3D scene using the edited images and then introduce a dynamic 4D Gaussian field to generate snowflakes, raindrops and fog in the scene. The attributes and dynamics of these particles are precisely controlled through physical-based modelling and simulation, ensuring realistic weather representation and flexible severity adjustments. Finally, we integrate the 4D Gaussian field with the 3D scene to render consistent and highly realistic weather effects. Experiments on multiple driving datasets demonstrate that WeatherEdit can generate diverse weather effects with controllable condition severity, highlighting its potential for autonomous driving simulation in adverse weather. See project page at https://jumponthemoon.github.io/w-edit/.
Introduction
@@
3D weather editing (3D-WE) aims to generate multi-weather conditions in 3D scenes based on user instructions. It is a crucial task in computer vision with applications in autonomous driving, augmented reality, and virtual scene synthesis (Almalioglu et al. 2022). While existing methods range from 2D-based image techniques (Valanarasu, Yasarla, and Patel 2022; Zhang et al. 2023) to 3D neural rendering (Li et al. 2023), developing a unified framework that enables multiple realistic weather effects and controllable severity remains a significant challenge. For instance, style transfer methods can modify weather backgrounds but are typically limited to a single effect. Diffusion models (Yang et al. 2024) enable multi-weather synthesis; however, they often lack fine-grained control and struggle with spatial and temporal consistency across multiple images (Rothmeier and Huber 2021). More recently, ClimateNeRF (Li et al. 2023) utilized Neural Radiance Fields (NeRF) to model snow accumulation and flooding in 3D space. However, this method is limited to static weather events and leaves the overall weather tone unchanged, lacking full-scene atmospheric realism.
Figure 1: Realistic, Controllable Weather Editing. Given input sunny weather images from multi-view cameras across multiple timesteps, WeatherEdit  can generate high-fidelity scenarios under diverse weather conditions (e.g., foggy, snowy, and rainy) with various severity levels (e.g., light, moderate, heavy) based on user instructions.
@@
    We demonstrate WeatherEdit’s effectiveness in generating realistic, consistent, and controllable weather effects in 3D driving scenes (shown in Figure 1), showcasing its applicability to real-world scenarios.
Related Work
@@
Existing methods can be categorized into 2D-based and 3D-based approaches. 2D-based methods utilize generative models (Zhu et al. 2017; Parmar et al. 2024) to synthesize weather effects in single images. (Hwang et al. 2022; Schmidt et al. 2019) use CycleGAN (Zhu et al. 2017) to edit weather conditions, while (Schmidt et al. ) applies GAN-based inpainting to generate flooding with water reflections. (Hahner et al. 2019) uses a graphical model to produce fog effects. Conventional 3D-based methods (Stomakhin et al. 2013; Zsolnai-Fehér 2022; Feldman and O’Brien 2002; Haas 2014) use graphical simulation to model weather in 3D space. More recently, (Li et al. 2023) has introduced neural rendering techniques for creating weather backgrounds but it is limited to post-weather event.
@@
Diffusion models (Ho, Jain, and Abbeel 2020; Meng et al. 2021; Rombach et al. 2021) have significantly advanced image editing, enabling flexible and controllable modifications. Methods such as ControlNet (Zhang, Rao, and Agrawala 2023) and T2I-Adapter (Mou et al. 2024) introduce trainable modules to guide the generation process, allowing for fine-grained control over diffusion-based outputs. InstructPix2Pix (Brooks, Holynski, and Efros 2023) enables interactive editing with textual instructions by trained on image-text pairs. TurboEdit (Deutch et al. 2024) employs a pseudo-guidance approach to enhance edit magnitude without introducing artifacts.
@@
Neural Radiance Fields (NeRF) (Mildenhall et al. 2020)and 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) are leading techniques for 3D scene reconstruction from multi-view images. NeRF encodes scenes implicitly with a neural network, while 3DGS represents them explicitly as 3D Gaussians, enabling faster rendering. Their versatility allows applications in dynamic scene reconstruction (Chen et al. 2025; Yan, Li, and Lee 2023), content creation (Liu et al. 2024; Lin et al. 2023), and artifact removal (Qian et al. 2025). While (Li et al. 2023) extends NeRF to model smog, snow, and floods for weather simulation but focuses only on static weather conditions, leaving dynamic weather phenomena unaddressed.
Methods
In order to generate realistic and controllable multi-weather scenes in 3D, we propose a novel framework, WeatherEdit, which follows a progressive 2D-to-4D transformation process. As illustrated in Figure 2, WeatherEdit  adopts a two-step weather editing pipeline: weather background editing and weather particle construction. The first step generates distortion-free, spatially and temporally aligned 2D images with edited backgrounds, while the second step dynamically models weather particles with various types and severity levels. The joint association of these two steps enables the flexible generation of multi-weather scenes.
Weather Background Editing
To support multiple editing styles, we develop a unified adapter to finetune the diffusion model and further enhance the results through semantically conditioned inputs. Additionally, we introduce a novel Temporal-View Attention mechanism to ensure consistency in the edited results.
All-in-one Adapter. Existing diffusion-based methods often suffer from scene deformations and excessive stylization, making them unsuitable for precise and realistic weather editing. Although (Parmar et al. 2024) achieves effective style editing by fine-tuning the models with multiple single-style LoRA (Hu et al. 2022) adapters (shown in Figure 3 (a)), it requires training a separate model for each style, resulting in increased computational cost and limited flexibility.
Figure 3: Single-style Adapters vs. All-in-One Adapter. (a) Utilizes separate adapters for each weather style (e.g., snow, rain, fog), while (b) employs a unified all-in-one adapter for efficient multi-style adaptation. Blue box highlights adapters to be trained.
To address this, we develop an all-in-one LoRA adapter that injects multiple weather styles into a single model (as shown in Figure 3 (b)), eliminating the need for training multiple ones. Let $W_{0}\in\mathbb{R}^{d\times k}$ be the pre-trained weights, $x$ the input image, and $i$ the weather style (e.g., snowy, rainy, foggy) specified by text prompt. Each style is associated with a low-rank matrix $L_{i}$. The adapted model’s forward pass is then computed as $h=W_{0}x+L_{i}x$.
@@ Methods
Temporal-View Consistency Alignment.
@@
After fine-tuning the model, multi-time and multi-view images are processed to achieve background scene editing. However, processing each image independently can lead to inconsistencies across time steps and viewpoints due to the inherent stochasticity of generative models. To address this, we propose a Temporal-View (TV-) attention mechanism (as shown in Figure 4) to enable our model to integrate information from adjacent frames and key viewpoints.
@@
$\text{S\_Attn}^{v}_{t}=\text{Softmax}\left(\frac{Q^{v}_{t}\cdot K^{v,T}_{t}}{\sqrt{d}}\right)V^{v}_{t},$ (2)
@@
To preserve temporal coherence, each frame $t$ queries its adjacent frames (t1t-1 and $t+1$), allowing it to integrate contextual information from both past and future frames. The temporal attention is defined as:
$\text{T\_Attn}^{v}_{t}=\text{Softmax}\left(\frac{Q^{v}_{t}\cdot K^{\mathcal{T}(t),T}_{v}}{\sqrt{d}}\right)V^{\mathcal{T}(t)}_{v}$, (3)
@@
Weather Particle Construction
Real-world weather is inherently regional, and we incorporate this into weather particle construction. Specifically, as shown in Figure 5, we propose a 4D Gaussian field, initialized with 3D Gaussians carrying physical attributes (e.g., color, quantity) to model raindrops, fog, and snowflakes through attribute modeling. To enhance realism, we apply dynamic simulation to create realistic weather effects.
@@
Experiments
Figure 6: Comparison with 2D-based Editing Methods. We show (a) original images with edited results from (b) ControlNet, (c) InstructPix2Pix, (d) TurboEdit, and (e) Ours. Existing methods often suffer from over-stylization, content removal, and inconsistencies in spatial and temporal coherence. In contrast, our method preserves scene integrity while ensuring multi-view and multi-frame consistency, producing realistic and temporally coherent weather effects across viewpoints and frames.
@@
Dataset. To develop the weather dataset for all-in-one adapter fine-tuning, we extract 1,237 image pairs from the BDD100K (Yu et al. 2020), MUSE (Brödermann et al. 2024), and ACDC (Sakaridis, Dai, and Van Gool 2021) datasets. Each pair contains a normal-weather image and an adverse-weather image, including snowy, rainy, and foggy conditions, along with a corresponding ground-truth semantic segmentation map. For weather editing, we selected eight scenes with multi-view image sequences from the Pandaset (Xiao et al. 2021), Waymo Open Dataset (Sun et al. 2020), nuScenes (Caesar et al. 2020), and KITTI-360 (Liao, Xie, and Geiger 2022) datasets.
@@
Implementation details. We implement the 2D editing model using CycleGAN-Turbo(Parmar et al. 2024), using a single RTX A6000, which takes 12 hours training and runs at 3.2 seconds per frame. The 3D scene reconstruction and 4D Gaussian field are based on OmniRe(Chen et al. 2025) and 3DGS(Kerbl et al. 2023), with both training(~1 hour) and rendering(~0.21s) performed on a single RTX 3090.
@@
Evaluation Metrics. To evaluate edited weather background quality, we use the cosine similarity of CLIP (Radford et al. 2021) image embeddings (CLIP-S) to measure content preservation and the directional CLIP similarity (Gal et al. 2022) (CLIP-DS) to assess alignment with text instructions. To assess consistency across temporal and spatial dimensions, we calculate the warp error to measure temporal coherence between adjacent frames and employ the Bhattacharyya distance (Bhattacharyya 1943) to quantify color distribution similarity across viewpoints.
Results
@@
Figure 8: Controllable Weather Effect and Severity in Driving Scenes. WeatherEdit can modify (a) original driving scenes from Pandaset, Nuscenes, and Waymo datasets by generating snowy, foggy, and rainy effects while preserving scene structure in (b). Additionally, it enables (c) controllable weather severity, adjusting precipitation intensity from light to heavy for realistic and flexible weather simulation.
Experimental Results
@@
We present both qualitative and quantitative comparisons of our weather background editing method against 2D-based approaches, including ControlNet (Zhang, Rao, and Agrawala 2023), InstructPix2Pix (Brooks, Holynski, and Efros 2023), and TurboEdit (Deutch et al. 2024), as well as 3D-based methods, such as 3D stylization (Li et al. 2018) and ClimateNeRF (Li et al. 2023). For dynamic weather effects, we apply qualitative analysis, i.e., controllable weather simulation, to demonstrate the results.
@@
Weather Background Editing Comparison. For the comparison of 2D-based methods, as shown in Figure 6, we observe that ControlNet tends to over-stylize the image, altering its original content. InstructPix2Pix produces a reasonable effect; however, the effect is inconsistent across adjacent frames and viewpoints. While TurboEdit accurately stylizes the image, it significantly modifies the scene content (e.g., turning pedestrians into water reflections and trees in a foggy scene). In contrast, our method achieves a realistic and coherent weather transformation while preserving scene integrity, as also indicated by the highest CLIP-DS and CLIP-S scores shown in Table 1.
ControlNet IP2P TurboEdit Ours
CLIP-DS Score ($\uparrow$) 0.291 0.295 0.257 0.302
CLIP-S Score ($\uparrow$) 0.594 0.691 0.676 0.750
Table 1: Comparison with 2D-based weather editing.
@@
CLIP-DS Score ($\uparrow$) 0.225 0.267 0.283
CLIP-S Score ($\uparrow$) 0.842 0.702 0.751
Table 2: Comparison with 3D-based weather editing.
For 3D-based methods, as shown in Figure 7, 3D stylization produces only a barely perceptible snowy effect on the vegetation, aligning with the highest CLIP-S score (shown in Table 2), which indicates minimal change in the scene. ClimateNeRF fails to modify the overall scene tone to accurately reflect the intended weather conditions. In contrast, our approach not only synthesizes a realistic snowy effect but also adjusts the scene tone to create a coherent and perceptually convincing snowy atmosphere, as also indicated by the highest CLIP-DS score in Table 2.
Controllable Weather Simulation. A key advantage of our method over existing approaches is its ability to generate dynamic weather effects while precisely controlling both the type and severity of weather conditions. Additionally, we can flexibly choose whether to introduce weather particles and, if included, adjust their intensity with fine granularity. This allows for a range of scenarios, from a static wet road surface without active rainfall to a heavy snowstorm with dense snowfall obscuring visibility. To visualize the results, we render edited weather backgrounds and modulate weather severity. As shown in Figure 8b, backgrounds are reconstructed from 2D editing outputs, with weather type determined by the selected adapters. We then apply the 4D Gaussian field to introduce particles (Figure 8c). By adjusting particle attributes such as color, opacity, shape, rotation, and velocity, we simulate snow, rain, and fog from light to heavy. This enables flexible, physically consistent scene manipulation for diverse weather conditions.
ACDC MUSE
w/o w w/o w
HRDA 46.3 58.7 (+12.4) 42.2 57.1 (+14.9)
MIC 50.0 59.0 (+9.0 ) 48.8 55.4 (+6.8)
Table 3: Comparison of segmentation results with and without our simulated weather data.
Benefits for Downstream Tasks. We validate the benefits of our simulation as data augmentation for a downstream semantic segmentation task on adverse weather dataset such as ACDC(Sakaridis, Dai, and Van Gool 2021), MUSE(Brödermann et al. 2024). We simulate 1273 frames of edited weather scenes and use HRDA(Hoyer, Dai, and Van Gool 2022) and MIC(Hoyer et al. 2023) for training. As shown in Table 3, our simulated data boosts the performance across both datasets and models, with mIoU gains up 14.9%. This demonstrates that our simulated scenes effectively enhance model robustness under adverse weather conditions, validating the utility of WeatherEdit for downstream tasks.
Ablation Studies
@@
Semantic Conditioned Input. Note that, although the proposed 2D background editing method relies on a semantic map as input (as shown in Figure 2), it only requires the semantic map generated under sunny weather conditions, which can be easily obtained using current open-source models. To evaluate the effectiveness of semantic-conditioned input, we compare editing results and FID scores with and without it. By computing the cross-attention map using “snowy” in the text prompt, we observe that semantic conditioning enhances the model’s response to sidewalks and trees, aligning weather effects more accurately with scene structures, as shown in Figure 11. Additionally, the model achieves a lower average FID value across different weather conditions compared to cases without semantic conditioning, demonstrating its effectiveness in capturing weather characteristics during training.
Temporal-View Attention. As shown in Figure 10, without TV-Attention, view 3 exhibits inconsistent editing effects, with a darker tone across two frames compared to view 2. After incorporating temporal-view attention, view 3 achieves more consistent editing, displaying a similar snowy effect between adjacent views and frames. For the qualitative analysis shown in Table 4, temporal attention reduces warp error, indicating improved temporal coherence, while spatial attention results in a lower Bhattacharyya distance, suggesting more consistent editing across viewpoints. By integrating both temporal and view attention, the full implementation achieves the lowest Bhattacharyya distance while maintaining a balanced reduction in warp error, ensuring globally consistent editing results.
Figure 9: Ablation on WeatherEdit modules.
@@
Self-attn +View +Temp Full
Warp-error ($\downarrow$) 0.043 0.042 0.039 0.041
Bhattacharyya dist ($\downarrow$) 0.272 0.253 0.261 0.245
@@
Modular Design for 4D Gaussian Field. As the cross-frame results in Figure 9 show, without 2D background editing (b), the scene lacks an overall weather tone. Randomly initializing 4D Gaussians for attribute modeling (c) produces colorful blobs instead of realistic weather. Without dynamic simulation (d), the particles remain relatively static in the camera view, lacking the natural falling motion. Without local field alignment (e), weather particles stay fixed at the initial position of the scene. As the camera moves forward, the weather particles gradually disappear instead of persisting throughout the scene. In contrast, when all modules are included, the full implementation (f) generates a realistic weather background with natural dynamics.
Conclusion
In this work, we introduce WeatherEdit, a novel 4D weather editing framework capable of generating realistic weather effects such as fog, rain, and snow, with fine-grained control over their severity. Our approach leverages an all-in-one adapter that enables a diffusion model to transform normal-weather images into diverse weather styles. To maintain consistency of edited results across multiple frames and viewpoints, we introduce an ordered Temporal-View (TV) Attention mechanism. Additionally, by designing a 4D Gaussian field, we model the attributes and dynamics of weather particles within a 3D scene, enhancing realism and enabling fine-grained control over weather severity. WeatherEdit demonstrates its potential in generating extreme weather conditions from normal scenes, providing a foundation for future research on evaluating the resilience of real-world applications such as autonomous driving.